OpenStudy (anonymous):

how to get vertex form of f(x)=4x^2+8x+2

OpenStudy (anonymous):

\[ f(x)=4x^2+8x+2\\ f(x) = 4x^2 + 8x +4-4+2\\ f(x) = (2 x+2)^2-2\\ \] Can you finish it now?

OpenStudy (anonymous):

eh I dont think so

OpenStudy (anonymous):

what x makes 2 x+2 =0?

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

f(-1)=-2

OpenStudy (anonymous):

vertex=(-1,-2)

OpenStudy (anonymous):

I need it in vertex form like f(x)=4(x+2)^2-6

OpenStudy (anonymous):

\[ f(x)=4x^2+8x+2\\ f(x) = 4x^2 + 8x +4-4+2\\ f(x) = (2 x+2)^2-2\\ f(x) = 4( x+1)^2-2\\ \]

OpenStudy (anonymous):

Did you get it now?

OpenStudy (anonymous):

ok why did you turn the positive 2 too a negative in the 3rd step

OpenStudy (anonymous):

-4+2=-2

OpenStudy (anonymous):

give me a sec

OpenStudy (anonymous):

the second step is whats getting me, why did you put +4-4+2

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

0=4-4 I added zero

OpenStudy (anonymous):

Have you heard about completing the square?

OpenStudy (anonymous):

thats what I was trying to find out previously, I have no clue on how to do it

OpenStudy (anonymous):

\[ 4 x^2 +8 x + 4=(2x+2)^2\\ (2x)^2 + 2(2x)2+ 2^2\\ a^2+ 2 a b + b^2=(a+b)^2 \] I have to add 4 to make it a perfect square and to not do any damage I take away -4

OpenStudy (anonymous):

Forget it i'm wasting time here not mine only but yours as well nevermind. Thanks for helping me. @eliassaab

OpenStudy (anonymous):

YW